how do u evaluate each expression for the replacement set{3,4,9} the question is a-2
Answer by jim_thompson5910(35256)

The replacement set is just the set of all the values you'll plug in for 'a'

Ex: Plug in the first number 3 from the replacement set to get: a-2=3-2=1


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website.

Jim